OpenStudy (anonymous):

solve y = 2x - 2. A) 1,3 B) 2,5 C) 2,3 D) 2,2 Help Explain

OpenStudy (anonymous):

Well the question is asking you to check and see. So if we were to take the first (x,y) coordinate and plug it in it'd be 3 = 2(1)-2 5 = 2(2)-2 3 = 2(3)-2 2 = 2(2)-2 So which statement is true?

OpenStudy (anonymous):

It would be 2=2(2)-2

OpenStudy (anonymous):

Yep because 2=2. Therefore since Left Side = Right Side you're solid

OpenStudy (anonymous):

Thank You so Much :)

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

so (2,2)

OpenStudy (anonymous):

Yep because that is the (x,y) coordinate

OpenStudy (anonymous):

Ok thanks again